Circular No. 5641

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

γ-RAY BURSTS C. A. Meegan, G. J. Fishman, R. B. Wilson, W. S. Paciesas, M. N.

Brock, J. M. Horack, G. N. Pendleton, and C. Kouveliotou report for the BATSE Science Team: “The BATSE experiment on the Compton Gamma Ray Observatory has detected 447 cosmic γ-ray bursts from 1991 Apr. 24 through 1992 Oct. 7. An improved algorithm for computing the directions to the bursts has been developed, which reduces systematic errors to ∼ 2^{o} (versus ∼ 6^{o} previously). So far, a total of 238 burst locations have been recomputed with the new algorithm. The distribution of 447 located events is isotropic within the statistical limits, as we have reported previously (IAUC 5358; Nature 355, 143; IAUC 5478). A measure of the dipole moment with respect to the Galactic Center is <cos θ> = 0.034 ± 0.027 (versus –0.014 for an isotropic distribution, including the effect of non-uniform sky coverage). A measure of the quadrupole moment with respect to the Galactic Plane is <sin^{2} b> = 0.316 ± 0.014 (versus 0.329 for an isotropic distribution, including the effect on non-uniform sky coverage). V/V_{max} is 0.324 ± 0.016. The first BATSE burst catalogue will be made publicly available through the GRO Science Support Center within the next few weeks. It will contain bursts detected between launch in 1991 April and 1992 Mar.

5. The catalogue may be accessed via SPAN (SET HOST GROSSC) or Telnet (grossc.gsfc.nasa.gov); the username is ‘gronews’.”

COMET SHOEMAKER-LEVY (1991a_{1})

Ephemeris from orbital elements on MPC 20602:

1992/93 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 15 12^{h}22.^{m}26 −36^{o}01.^{′}6 2.433 1.654 30.^{o}6 17.^{o}9 11.6
25 12 23.46 −40 04.9 2.509 1.787 35.0 18.6 12.0
Nov. 4 12 23.52 −44 09.7 2.564 1.920 40.4 19.6 12.4
14 12 21.76 −48 18.2 2.601 2.051 46.6 20.5 12.7
24 12 17.20 −52 31.1 2.626 2.181 53.1 21.2 13.0
Dec. 4 12 08.37 −56 46.1 2.641 2.310 59.9 21.7 13.2
14 11 53.11 −60 56.8 2.653 2.437 66.7 21.8 13.5
24 11 28.36 −64 50.5 2.666 2.562 73.3 21.6 13.7
Jan. 3 10 50.47 −68 04.8 2.685 2.686 79.5 21.1 13.9
13 9 58.1 −70 07.5 2.715 2.809 85.1 20.4 14.2

1992 October 21 Daniel W. E. Green